# A complete classification of the zero-dimensional homogeneous spaces under determinacy

### v2

@@ 9.2. Homogeneous spaces have good complexity
Proof.
Assume without loss of generality that $X$ is a subspace of $\omega^{\omega}$. Let $A$ be a non-empty open $\mathbf{\Gamma}$ subspace of $X$. Define
$\mathcal{U}=\{h[A]:h\text{ is a homeomorphism of }X\}.$
Notice that $\mathcal{U}$ is a cover of $X$ because $X$ is homogeneous and $A$ is non-empty. Let $\{A_{n}:n\in\omega\}$ be a countable subcover of $\mathcal{U}$, and observe that each $A_{n}\in\mathbf{\Gamma}$ by Proposition 9.3. Fix $U_{n}\in\mathbf{\Sigma}^{0}_{1}(\omega^{\omega})$ for $n\in\omega$ such that each $U_{n}\cap X=A_{n}$. Set
$V_{-1}=\omega^{\omega}\setminus\bigcup_{k\in\omega}U_{k},\text{ and }V_{n}=U_{n}\setminus\bigcup_{k<n}U_{k}$
for $n\in\omega$. Notice that $V_{n}\in\mathbf{\Delta}^{0}_{2}(\omega^{\omega})$ whenever $-1\leq n<\omega$, and that $V_{m}\cap V_{n}=\varnothing$ whenever $-1\leq m,n<\omega$ and $m\neq n$. Furthermore, it is easy to check that
$X=\bigcup_{-1\leq n<\omega}(A_{n}\cap V_{n}),$
where $A_{-1}=\varnothing$. In conclusion, we have shown that $X\in\mathsf{PU}_{1}(\mathbf{\Gamma})$. Since $\ell(\mathbf{\Gamma})\geq 1$, it follows that $X\in\mathbf{\Gamma}$. ∎
@@
Proof.
Start by fixing an embedding $i:X\longrightarrow 2^{\omega}$. Notice that $X$ must be crowded, otherwise $X$ would be discrete by homogeneity. Therefore, by taking the closure of $i[X]$ in $2^{\omega}$ and applying [Ke, Theorem 7.4], we can assume without loss of generality that $i$ is a dense embedding. Observe that $i[X]$ must be codense in $2^{\omega}$, otherwise $X$ would be locally compact by homogeneity. So we can fix a countable dense subset $D$ of $2^{\omega}$ such that $D\cap i[X]=\varnothing$. Since $2^{\omega}\setminus D\approx\omega^{\omega}$ by [Ke, Theorem 7.7], this means that there exists a dense embedding $j:X\longrightarrow\omega^{\omega}$. Furthermore, since $X$ is a $\mathbf{\Sigma}$ space and $\mathbf{\Sigma}$ is a nice topological pointclass, we must have $j[X]\in\mathbf{\Sigma}(\omega^{\omega})$. For the sake of simplicity, from now on, we will actually assume that $X\in\mathbf{\Sigma}(\omega^{\omega})$ and that $X$ is dense in $\omega^{\omega}$.
By Theorem 3.4, it is possible to fix a $\subseteq$-minimal $\mathbf{\Gamma}\in\mathsf{NSD}_{\mathbf{\Sigma}}(\omega^{\omega})$ such that $X\cap U\in\mathbf{\Gamma}\cup\widecheck{\mathbf{\Gamma}}$ for some non-empty $U\in\mathbf{\Delta}^{0}_{1}(\omega^{\omega})$. Fix a non-empty $U\in\mathbf{\Delta}^{0}_{1}(\omega^{\omega})$ such that $X\cap U\in\mathbf{\Gamma}\cup\widecheck{\mathbf{\Gamma}}$. We will assume that $X\cap U\in\mathbf{\Gamma}$, since the case $X\cap U\in\widecheck{\mathbf{\Gamma}}$ is similar. Clearly, the following two claims will conclude the proof.
Claim 1. $\mathbf{\Gamma}$ is a good Wadge class.
Proof. By applying Lemma 9.6 to the good Wadge class $\mathsf{D}_{\omega}(\mathbf{\Sigma}^{0}_{2}(\omega^{\omega}))$ and its dual, one sees that $X\cap U\notin\Delta(\mathsf{D}_{\omega}(\mathbf{\Sigma}^{0}_{2}(\omega^{\omega})))$, otherwise the assumption that $X$ is not a $\Delta(\mathsf{D}_{\omega}(\mathbf{\Sigma}^{0}_{2}))$ space would be contradicted by Proposition 9.3. By Lemma 3.3, it follows that $\Delta(\mathsf{D}_{\omega}(\mathbf{\Sigma}^{0}_{2}(\omega^{\omega})))\subseteq\mathbf{\Gamma}$. It remains to show that $\ell(\mathbf{\Gamma})\geq 1$. Assume, in order to get a contradiction, that $\ell(\mathbf{\Gamma})=0$. Observe that this implies $\ell(\mathbf{\Gamma}(U))=0$ by Theorem 3.20. Also notice that $X$ must be codense in $\omega^{\omega}$, otherwise $X$ would be Polish by homogeneity. Since $X\cap U\in\mathbf{\Gamma}(U)$ by Lemma 3.12, it follows from Lemma 3.22 that there exists a non-empty $V\in\mathbf{\Delta}^{0}_{1}(U)$ and $\mathbf{\Lambda}\in\mathsf{NSD}_{\mathbf{\Sigma}}(\omega^{\omega})$ such that $\mathbf{\Lambda}(U)\subsetneq\mathbf{\Gamma}(U)$ and $X\cap V\in\mathbf{\Lambda}(U)$. Observe that $\mathbf{\Lambda}\subsetneq\mathbf{\Gamma}$ by Theorem 3.14. Therefore, in order to contradict the minimality of $\mathbf{\Gamma}$, it will be enough to show that $X\cap V\in\mathbf{\Lambda}$. By Lemma 3.12, there exists $A\in\mathbf{\Lambda}$ such that $A\cap U=X\cap V$. Notice that $\mathbf{\Lambda}\neq\{\omega^{\omega}\}$ because $X$ is codense in $\omega^{\omega}$. Therefore $X\cap V=A\cap U\in\mathbf{\Lambda}$ by Lemma 3.7. $\blacksquare$
Claim 2. $X\!\!\downarrow\,=\mathbf{\Gamma}$.
Proof. By Claim 1, it is possible to apply Lemma 9.6, which yields $X\in\mathbf{\Gamma}$ by Proposition 9.3. This means that $X\!\!\downarrow\,\subseteq\mathbf{\Gamma}$. To conclude the proof, it will be enough to show that $X$ is non-selfdual, since the desired result will then follow from the minimality of $\mathbf{\Gamma}$. Assume, in order to get a contradiction, that $X$ is selfdual. Then, by Theorem 3.6, there exist a non-empty $V\in\mathbf{\Delta}^{0}_{1}(\omega^{\omega})$ and a non-selfdual $A<X$ in $\omega^{\omega}$ such that $A\cap V=X\cap V$. By the same argument used at the end of the proof of Claim 1, this contradicts the minimality of $\mathbf{\Gamma}$. $\blacksquare$
9.3. Uniqueness
In this section we will show that, up to homeomorphism, there exists at most one zero-dimensional homogeneous space of a given exact complexity and Baire category. In other words, for every good Wadge class $\mathbf{\Gamma}$ in $\omega^{\omega}$, there exist at most two zero-dimensional homogeneous spaces of exact complexity $\mathbf{\Gamma}$: one meager, and one Baire.222 Recall that, by Proposition 2.21, every homogeneous space is either meager or Baire. This is the uniqueness part of the classification. Theorem 9.9 is essentially the same as [CMM1, Theorem 15.2] (the difference is that the statement given here is level-by-level, and it uses the notions of complexity introduced in Section 9.1).
@@ 9.3. Uniqueness
We decided to state this result in the following form (that is, in terms of embeddings into $2^{\omega}$) for two reasons. First, this will allow us to derive Corollary 9.10, which will be useful in the proofs of Corollary 9.17 and Theorem 9.19. Second, Theorem 9.9 shows that, given a zero-dimensional homogeneous space $X$ of sufficiently high complexity, there is essentially a unique way to embed $X$ into $2^{\omega}$ densely, which seems to be of independent interest.
We begin with a simple technical result, which will allow us to apply Steel’s Theorem 8.9. Lemma 9.8 first appeared as [CMM1, Lemma 13.2].
Lemma 9.8 (Carroy, Medini, Müller).
Let $\mathbf{\Sigma}$ be a nice topological pointclass, and assume that $\mathsf{Det}(\mathbf{\Sigma}(\omega^{\omega}))$ holds. If $\mathbf{\Gamma}\subseteq\mathbf{\Sigma}(2^{\omega})$ is a good Wadge class in $2^{\omega}$ then $\mathbf{\Gamma}$ is reasonably closed.
Proof.
Let $\mathbf{\Gamma}(2^{\omega})\in\mathsf{NSD}_{\mathbf{\Sigma}}(2^{\omega})$ be a good Wadge class, where $\mathbf{\Gamma}\in\mathsf{NSD}_{\mathbf{\Sigma}}(\omega^{\omega})$. Set $Z=2^{\omega}\setminus(Q_{0}\cup Q_{1})$, and let $\phi:Z\longrightarrow 2^{\omega}$ be the function defined in Section 8.3. Pick $A\in\mathbf{\Gamma}(2^{\omega})$. Notice that $\phi^{-1}[A]\in\mathbf{\Gamma}(Z)$ by Lemma 3.11. Therefore, by Lemma 3.12, there exists $B\in\mathbf{\Gamma}(2^{\omega})$ such that $B\cap Z=\phi^{-1}[A]$. But $\mathbf{\Gamma}(2^{\omega})$ is good, hence $\phi^{-1}[A]\in\mathbf{\Gamma}(2^{\omega})$ by Corollary 7.13. Finally, a further application of Corollary 7.13 shows that $\phi^{-1}[A]\cup Q_{0}\in\mathbf{\Gamma}(2^{\omega})$, as desired. ∎
@@
Proof.
Let $i:X\longrightarrow 2^{\omega}$ and $j:Y\longrightarrow 2^{\omega}$ be dense embeddings. Observe that $\mathbf{\Gamma}(2^{\omega})$ is a good Wadge class by Lemma 7.10, hence it is reasonably closed by Lemma 9.8. Therefore, by Theorem 8.9, it will be enough to prove the following three claims.
Claim 1. $(i[X]\cap\mathsf{N}_{s})\!\!\downarrow\,=\mathbf{\Gamma}(2^{\omega})$ for every $s\in 2^{<\omega}$.
Proof of Claim 1. Pick $s\in 2^{<\omega}$. Since $i[X]$ is clearly a $\mathbf{\Gamma}$ space, an application of Proposition 9.5 shows that $i[X]\cap\mathsf{N}_{s}$ is a $\mathbf{\Gamma}$ space. Furthermore, since $i[X]$ is not a $\widecheck{\mathbf{\Gamma}}$ space by Corollary 9.4, it follows from Lemma 9.6 that $i[X]\cap\mathsf{N}_{s}$ is a not a $\widecheck{\mathbf{\Gamma}}$ space. A further application of Corollary 9.4 concludes the proof. $\blacksquare$
Claim 2. $(j[Y]\cap\mathsf{N}_{s})\!\!\downarrow\,=\mathbf{\Gamma}(2^{\omega})$ for every $s\in 2^{<\omega}$.
Proof of Claim 2. Proceed as in the proof of Claim 1. $\blacksquare$
Claim 3. $i[X]$ and $j[Y]$ are either both meager in $2^{\omega}$ or both comeager in $2^{\omega}$.
Proof of Claim 3. If $X$ and $Y$ are both meager spaces, then $i[X]$ and $j[Y]$ will both be meager subsets of $2^{\omega}$ by density. If $X$ and $Y$ are both Baire spaces, then $i[X]$ and $j[Y]$ will both be comeager subsets of $2^{\omega}$ by Corollary 2.8. $\blacksquare$
@@ 9.4. Existence: the positive results
In this section, we will show that certain assumptions on the level, type, and separation property of a good Wadge class $\mathbf{\Gamma}$ permit the construction of a zero-dimensional homogeneous space of exact complexity $\mathbf{\Gamma}$. Our method of proof will naturally yield semifilters (recall that these are homogeneous by Corollary 2.18).
@@
  • $\widehat{S}=\{z\in Z:z\subseteq z^{\prime}\text{ for some }z^{\prime}\in S\}$ whenever $S\subseteq K$,
@@
=$\mathcal{R}=\{y\cup e:y\in Y\text{ and }e\in\mathsf{Fin}(\Omega)\},$
@@
Observe that each $A_{n}\in\mathbf{\Gamma}(Z)$ by Proposition 9.3, and that each $A_{n}$ is a closed subset of $\mathcal{R}$. Furthermore, it is clear that $\mathcal{R}=\bigcup_{n\in\omega}A_{n}$. Since $\mathbf{\Gamma}(Z)$ is closed under $\mathsf{SU}_{1}$ by Lemma 7.4, it follows from Lemma 9.12 that $\mathcal{R}\in\mathbf{\Gamma}(Z)$. To complete the proof, simply observe that $\mathcal{R}\cap K=X$, then apply Lemmas 7.11 and 9.11. $\blacksquare$
@@ 9.5. Existence: the negative results
Next, we will prove that the results of Section 9.4 are sharp. More precisely, as Theorem 9.16 and Corollary 9.17 show, there are no other zero-dimensional homogeneous spaces whose exact complexity is a good Wadge class. Theorem 9.16 is based on [vE3, Lemma 4.3.7].
@@
Let $X$ be a zero-dimensional homogeneous space of exact complexity $\mathbf{\Gamma}$. Assume without loss of generality that $X$ is a dense subspace of $2^{\omega}$. Set $Y=2^{\omega}\setminus X$, and observe that $Y$ is homogeneous by Corollaries 2.16 and 9.10. Furthermore, using Corollary 9.4, one sees that $Y$ has exact complexity $\widecheck{\mathbf{\Gamma}}$. Since $\widecheck{\mathbf{\Gamma}}$ does not have the separation property by Theorem 5.8, it follows from Theorem 9.16 that $Y$ is a meager space, hence a meager subset of $2^{\omega}$ by density. In conclusion, being a comeager subspace of $2^{\omega}$, the space $X$ is Baire. ∎
@@ 9.6. The classification of the zero-dimensional homogeneous spaces
As the following theorem shows, the previous three sections can be combined to yield a complete classification of the zero-dimensional homogeneous spaces of sufficiently high complexity under $\mathsf{AD}$. This classification is complete because, by Theorem 9.7, a zero-dimensional homogeneous space that is not $\mathbf{\Delta}(\mathsf{D}_{\omega}(\mathbf{\Sigma}^{0}_{2}))$ has exact complexity $\mathbf{\Gamma}$ for some good Wadge class $\mathbf{\Gamma}$. The case $\mathbf{\Sigma}=\mathsf{Borel}$ is due to van Engelen (see [vE3, Theorem 4.4.4]), although his results are expressed using the more complicated language of [Lo1].
@@
The existence claims follow from Corollaries 9.14 and 9.15, since every semifilter is homogeneous by Corollary 2.18. The constraints on Baire category follow from Theorem 9.16 and Corollary 9.17. Finally, the claims about uniqueness up to homeomorphism follow from Proposition 2.21 and Theorem 9.9. ∎
@@ 9.7. A purely topological characterization of semifilters
The following “self-strengthening” result is a natural byproduct of our classification. In particular, the equivalence $(1)\leftrightarrow(3)$ can be viewed as a purely topological characterization of semifilters. In Section 10.3, we will give an analogous characterization of filters. Theorem 9.19 simultaneously generalizes [vE3, Corollary 4.4.6] (which gives $(1)\leftrightarrow(2)$ in the case $\mathbf{\Sigma}=\mathsf{Borel}$), [Me4, Theorem 1.1] (which gives $(1)\leftrightarrow(3)$ in the case $\mathbf{\Sigma}=\mathsf{Borel}$), and [CMM1, Theorem 1.1] (which gives $(1)\leftrightarrow(2)$ in the case $\mathbf{\Sigma}=\mathcal{P}$). Furthermore, it answers [Me4, Question 13.4] in the affirmative.
@@
Chapter 10 Filters
In this chapter, we will build on the results of the previous one to classify filters up to homeomorphism (see Theorem 10.4). As a byproduct, we will give natural topological properties (originally obtained by van Engelen in [vE5] for the Borel case) that describe exactly when a zero-dimensional space is homeomorphic to a filter (see Theorem 10.5). In addition to closure under $\mathsf{SU}_{1}$ and the properties of good Wadge classes, we will make use of the notion of closure under squares. We will conclude the chapter by discussing the relationship between filters and topological groups.
@@ 10.1. Closure under squares
The proof of the implication $(3)\rightarrow(4)$ is similar to the one that we have just given, so we will omit it. The implications $(2)\rightarrow(1)$ and $(4)\rightarrow(3)\rightarrow(1)$ are trivial. To conclude the proof, we will show that $(1)\rightarrow(3)$. So fix an uncountable zero-dimensional Polish space $Z$ such that $\mathbf{\Gamma}(Z)$ is closed under squares. Let $2\leq n<\omega$ and $A\in\mathbf{\Gamma}(Z)$. Assume without loss of generality that $A\neq\varnothing$, and fix $x\in A$. Pick $k\in\omega$ such that $2^{k}\geq n$. Notice that $A^{2^{k}}\in\mathbf{\Gamma}(Z^{2^{k}})$ because $\mathbf{\Gamma}(Z)$ is closed under squares. Set
$W=\left\{z\in Z^{2^{k}}:z(i)=x\text{ whenever }n\leq i<2^{k}\right\},$
and observe that $A^{2^{k}}\cap W\in\mathbf{\Gamma}(W)$ by Lemma 3.12. Define $h:Z^{n}\longrightarrow W$ by setting $h(z_{0},\ldots,z_{n-1})=(z_{0},\ldots,z_{n-1},x,\ldots,x)$, and observe that $h$ is a homeomorphism. Therefore
$A^{n}=h^{-1}\left[A^{2^{k}}\cap W\right]\in\mathbf{\Gamma}(Z^{n})$
@@ 10.2. The topological classification of filters
In this section, we will give the promised classification of filters up to homeomorphism (see Theorem 10.4). Naturally, this classification will build on the results that we already have. The new existence result that we will need is Theorem 10.3, whose proof is taken almost verbatim from the proof of [vE5, Lemma 3.3]. However, as in the statement of Theorem 9.13, we have isolated more clearly the required closure properties. More importantly, in the proof of Theorem 10.4, we will use the results of Section 7.5.
@@
  • $\widehat{S}=\{z\in Z:z\subseteq z^{\prime}\text{ for some }z^{\prime}\in S\}$ whenever $S\subseteq K$,
@@
=$\mathcal{I}=\{y\cup e:y\in Y^{\prime}\text{ and }e\in\mathsf{Fin}(\Omega)\}.$
@@
Using Claim 4, Lemma 3.7 and Proposition 9.3, one sees that each $A_{k}\in\mathbf{\Gamma}(Z)$. Furthermore, each $A_{k}\in\mathbf{\Pi}^{0}_{1}(\mathcal{I})$ by Claim 3. Since $\mathcal{I}=\bigcup_{k\in\omega}A_{k}$, it follows from Lemmas 7.4 and 9.12 that $\mathcal{I}\in\mathbf{\Gamma}(Z)$. To complete the proof, simply observe that $\mathcal{I}\cap K=X$, then apply Lemmas 7.11 and 9.11. $\blacksquare$
@@
In order to prove the implication $(1)\rightarrow(2)$, let $\mathcal{F}$ be a filter of exact complexity $\mathbf{\Gamma}$. Notice that $\mathcal{F}$ is meager by Corollary 2.12. Therefore, by Theorem 9.18, it will be enough to show that $\mathbf{\Gamma}$ is closed under squares. By Proposition 10.2, it will be enough to show that $\mathbf{\Gamma}(2^{\omega})$ is closed under squares. So pick $A\in\mathbf{\Gamma}(2^{\omega})$, and observe that $A\leq\mathcal{F}$ in $2^{\omega}$ by Corollary 9.4. This trivially implies that $A\times A\leq\mathcal{F}\times\mathcal{F}$ in $2^{\omega}\times 2^{\omega}$. Since ×$\mathcal{F}\times\mathcal{F}\in\mathbf{\Gamma}(2^{\omega}\times 2^{\omega})$ by Theorem 2.10 and Proposition 9.3, it follows that $A\times A\in\mathbf{\Gamma}(2^{\omega}\times 2^{\omega})$, as desired. Now assume that $\mathbf{\Gamma}$ satisfies condition $(2)$. As in the proof of Corollary 9.14, one sees that $\mathbf{\Gamma}$ is closed under $\mathsf{SU}_{1}$. Therefore, condition $(1)$ holds by Theorem 10.3. ∎

### v3

@@ 6.2. Uniqueness
We finally come to the uniqueness result, which is essentially the same as [CMM1, Theorem 15.2]. We decided to state this result in the following form (that is, in terms of embeddings into $2^{\omega}$) for two reasons. First, this will allow us to derive Corollary 6.9, which will be useful in the proofs of Corollary 6.16 and Theorem 6.18. Second, Theorem 6.8 shows that, given a zero-dimensional homogeneous space $X$ of sufficiently high complexity, there is essentially a unique way to embed $X$ into $2^{\omega}$ densely, which seems to be of independent interest.
@@ 6.3. Existence: the positive results
In this subsection, we will show that certain assumptions on the level, type, and separation property of a good Wadge class $\mathbf{\Gamma}$ permit the construction of a zero-dimensional homogeneous space of exact complexity $\mathbf{\Gamma}$. Our method of proof will naturally yield semifilters (recall that these are homogeneous by Corollary 2.17).
@@
[…]
@@
$\mathcal{R}=\{y\cup e:y\in Y\text{ and }e\in\mathsf{Fin}(\Omega)\},$
@@
Observe that each $A_{n}\in\mathbf{\Gamma}(Z)$ by Proposition 6.3, and that each $A_{n}$ is a closed subset of $\mathcal{R}$. Furthermore, it is clear that $\mathcal{R}=\bigcup_{n\in\omega}A_{n}$. Since $\mathbf{\Gamma}(Z)$ is closed under $\mathsf{SU}_{1}$ by Lemma 5.4, it follows from Lemma 6.11 that $\mathcal{R}\in\mathbf{\Gamma}(Z)$. To complete the proof, simply observe that $\mathcal{R}\cap K=X$, then apply Lemmas 5.11 and 6.10. $\blacksquare$
@@ 6.4. Existence: the negative results
Next, we will prove that the results of §6.3 are sharp. More precisely, as Theorem 6.15 and Corollary 6.16 show, there are no other zero-dimensional homogeneous spaces whose exact complexity is a good Wadge class. Theorem 6.15 is based on [vE3, Lemma 4.3.7].
@@
Let $X$ be a zero-dimensional homogeneous space of exact complexity $\mathbf{\Gamma}$. Assume without loss of generality that $X$ is a dense subspace of $2^{\omega}$. Set $Y=2^{\omega}\setminus X$, and observe that $Y$ is homogeneous by Corollaries 2.15 and 6.9. Furthermore, using Corollary 6.4, one sees that $Y$ has exact complexity $\widecheck{\mathbf{\Gamma}}$. Since $\widecheck{\mathbf{\Gamma}}$ does not have the separation property by Theorem 3.50, it follows from Theorem 6.15 that $Y$ is a meager space, hence a meager subset of $2^{\omega}$ by density. In conclusion, being a comeager subspace of $2^{\omega}$, the space $X$ is Baire. ∎
@@ 6.5. The classification of the zero-dimensional homogeneous spaces
As the following theorem shows, the previous three subsections can be combined to yield a complete classification of the zero-dimensional homogeneous spaces of sufficiently high complexity under $\mathsf{AD}$. This classification is complete because, by Theorem 6.7, a zero-dimensional homogeneous space that is not $\mathbf{\Delta}(\mathsf{D}_{\omega}(\mathbf{\Sigma}^{0}_{2}))$ has exact complexity $\mathbf{\Gamma}$ for some good Wadge class $\mathbf{\Gamma}$. The case $\mathbf{\Sigma}=\mathsf{Borel}$ is due to van Engelen (see [vE3, Theorem 4.4.4]), although his results are expressed using the more complicated language of [Lo1].
@@
The existence claims follow from Corollaries 6.13 and 6.14, since every semifilter is homogeneous by Corollary 2.17. The constraints on Baire category follow from Theorem 6.15 and Corollary 6.16. Finally, the claims about uniqueness up to homeomorphism follow from Proposition 2.20 and Theorem 6.8. ∎
@@ 6.6. A purely topological characterization of semifilters
The following “self-strengthening” result is a natural byproduct of our classification. In particular, the equivalence $(1)\leftrightarrow(3)$ can be viewed as a purely topological characterization of semifilters. In §7.3, we will give an analogous characterization of filters. Theorem 6.18 simultaneously generalizes [vE3, Corollary 4.4.6] (which gives $(1)\leftrightarrow(2)$ in the case $\mathbf{\Sigma}=\mathsf{Borel}$), [Me4, Theorem 1.1] (which gives $(1)\leftrightarrow(3)$ in the case $\mathbf{\Sigma}=\mathsf{Borel}$), and [CMM1, Theorem 1.1] (which gives $(1)\leftrightarrow(2)$ in the case $\mathbf{\Sigma}=\mathcal{P}$). Furthermore, it answers [Me4, Question 13.4] in the affirmative.
@@
7. Filters
@@ 7.1. Closure under squares
The proof of the implication $(3)\rightarrow(4)$ is similar to the one that we have just given, so we will omit it. The implications $(2)\rightarrow(1)$ and $(4)\rightarrow(3)\rightarrow(2)$ are trivial. To conclude the proof, we will show that $(2)\rightarrow(4)$. So assume that $(2)$ holds. Pick a zero-dimensional Polish space $Z$, pick $A\in\mathbf{\Gamma}(Z)$ and $2\leq n<\omega$. Assume without loss of generality that $A\neq\varnothing$, and fix $x\in A$. Using $(2)$, it is easy to realize that $A^{m}\in\mathbf{\Gamma}(Z^{m})$ whenever $m=2^{k}$ for some $k\in\omega$. Now fix $k\in\omega$ such that $2^{k}\geq n$, then set $m=2^{k}$. Define
$W=\{z\in Z^{m}:z(i)=x\text{ whenever }n\leq i<m\},$
and observe that $A^{m}\cap W\in\mathbf{\Gamma}(W)$ by Lemma 3.10. Define $h:Z^{n}\longrightarrow W$ by setting $h(z_{0},\ldots,z_{n-1})=(z_{0},\ldots,z_{n-1},x,\ldots,x)$, and observe that $h$ is a homeomorphism. It follows that
$A^{n}=h^{-1}[A^{m}\cap W]\in\mathbf{\Gamma}(Z^{n})$
@@ 7.2. The topological classification of filters
In this subsection, we will give the promised classification of filters up to homeomorphism (see Theorem 7.4). Naturally, this classification will build on our previous results. The new existence result that we will need is Theorem 7.3, whose proof is taken almost verbatim from the proof of [vE5, Lemma 3.3]. However, as in the statement of Theorem 6.12, we have isolated more clearly the required closure properties. More importantly, in the proof of Theorem 7.4, we will use the results of §5.5.
@@
[…]
@@
$\mathcal{I}=\{y\cup e:y\in Y^{\prime}\text{ and }e\in\mathsf{Fin}(\Omega)\}.$
@@
Using Claim 4, Lemma 3.5 and Proposition 6.3, one sees that each $A_{k}\in\mathbf{\Gamma}(Z)$. Furthermore, each $A_{k}\in\mathbf{\Pi}^{0}_{1}(\mathcal{I})$ by Claim 3. Since $\mathcal{I}=\bigcup_{k\in\omega}A_{k}$, it follows from Lemmas 5.4 and 6.11 that $\mathcal{I}\in\mathbf{\Gamma}(Z)$. To complete the proof, simply observe that $\mathcal{I}\cap K=X$, then apply Lemmas 5.11 and 6.10. $\blacksquare$
@@
In order to prove the implication $(1)\rightarrow(2)$, let $\mathcal{F}$ be a filter of exact complexity $\mathbf{\Gamma}$. Notice that $\mathcal{F}$ is meager by Corollary 2.11. Therefore, by Theorem 6.17, it will be enough to show that $\mathbf{\Gamma}$ is closed under squares. By Proposition 7.2, it will be enough to show that $\mathbf{\Gamma}(2^{\omega})$ is closed under squares. So pick $A\in\mathbf{\Gamma}(2^{\omega})$, and observe that $A\leq\mathcal{F}$ in $2^{\omega}$ by Corollary 6.4. This trivially implies that $A\times A\leq\mathcal{F}\times\mathcal{F}$ in $2^{\omega}\times 2^{\omega}$. Since $\mathcal{F}\times\mathcal{F}\in\mathbf{\Gamma}(2^{\omega}\times 2^{\omega})$ by Theorem 2.9 and Proposition 6.3, it follows that $A\times A\in\mathbf{\Gamma}(2^{\omega}\times 2^{\omega})$, as desired. Now assume that $\mathbf{\Gamma}$ satisfies condition $(2)$. As in the proof of Corollary 6.13, one sees that $\mathbf{\Gamma}$ is closed under $\mathsf{SU}_{1}$. Therefore, condition $(1)$ holds by Theorem 7.3. ∎